OpenStudy (marcelie):

can someone give me an example of direct comparison test and limit comparison test

OpenStudy (marcelie):

@zepdrix

zepdrix (zepdrix):

\[\large\rm \sum_{n=1}^{\infty}\frac{1}{n^3+7}\]Hmm any ideas for this one?

OpenStudy (marcelie):

hmm i was thinking it might me like this 

zepdrix (zepdrix):

Comparing it to that? Ok seems like a good idea. Are the terms of \(\large\rm \frac{1}{n^3+7}\) going to be larger or smaller than the terms of \(\large\rm \frac{1}{n^3}\) ?

OpenStudy (marcelie):